OpenStudy (angiepangie0726):

Determine the zeros of x2 + 19x = 6x − 30.

OpenStudy (angiepangie0726):

@pina_colada123

OpenStudy (angiepangie0726):

@Britbrat1997

OpenStudy (misty1212):

HI!!

OpenStudy (angiepangie0726):

Hello :)

OpenStudy (angiepangie0726):

Do you think you can help me out?

OpenStudy (misty1212):

subtract \(6x\) from both sides and add 30 that way it will be set equal to zero

OpenStudy (angiepangie0726):

Ok! I will do that now

OpenStudy (angiepangie0726):

x^2+13x+30=0

OpenStudy (misty1212):

once you do that, it will factor nicely

OpenStudy (misty1212):

yeah like that now factor it, easy for this one

OpenStudy (angiepangie0726):

I am honestly the worst at factoring so please forgive me if I am wrong. My teacher had taught me a trick and she had said that if you take the middle number and divide by 2 that will be your number. I got 6.5 but I am not sure if that is correct or not

OpenStudy (misty1212):

no not hardly think of two numbers that multiply to give 30 and add to give 13

OpenStudy (angiepangie0726):

Oh 10 and 3 :)

OpenStudy (misty1212):

don't think too hard, this one you get right away right \[(x+3)(x+10)=0\]set each factor equal to zero and solve i am sure you can do it gtg

OpenStudy (angiepangie0726):

soooo x^2+3x+10x+30=0

OpenStudy (angiepangie0726):

ok thank you so much!

OpenStudy (angiepangie0726):

would it be -3,-10?